A firm has Sh 3,000,000 debt at 5% interest in its capital structureand the expected annual operation income is sh 1,000,000 and equity capitalization rate is 10%. Calculate the value of the firm and overall cost of capital.
1
Expert's answer
2019-04-03T09:04:32-0400

"Firm value = Debt value + Equity value"

"Firm value = 3,000,000 + 1,000,000\times0.1 = 3,100,000."

"Cost of capital = \frac{3,000,000\times0.05}{100,000} = 1.5" or 150%.